REFERENCES

[1]. J.F. Adams, Stable homotopy and generalised cohomology, University of Chicago Press,

Chicago, 1974. [2]. J.F. Adams, Infinite Loop Spaces, Princeton University Press, Princeton, 1978. [3]. M. Artin and B. Mazur, Étale homotopy, Lecture Notes in Math., Vol. 100, Springer,

Berlin-Heidelberg-New York, 1972. [4]. M. Barr and C. Wells, Toposes, Triples and Theories, Springer-Verlag, New York, 1985. [5]. M. Barratt, V.K.A.M. Guggenheim and J.C. Moore, On semi-simplicial fibre bundles,

Amer. J. Math. 81 (1959), 639–657. [6]. M.G. Barratt and S. Priddy, On the homology of non-connected monoids and their as-

sociated groups, Comment. Math. Helv. 47 (1972), 1–14. [7]. A.K. Bousfield, A second quadrant homotopy spectral sequence, Trans. AMS 177 (1973),

305–318. [8]. A.K. Bousfield, The localization of spaces with respect to homology, Topology 14 (1975),

133–150. [9]. A.K. Bousfield, The localization of spectra with respect to homology, Topology 18 (1979),

257–281. [10]. A.K. Bousfield, Homotopy spectral sequences and obstructions, Isr. J. Math. 66 (1989),

54-104. [11]. A.K. Bousfield, Localization and periodicity in unstable homotopy theory, J. AMS 7

(1994), 831–873. [12]. A.K. Bousfield and E.M. Friedlander, Homotopy theory of Γ-spaces, spectra, and bisim-

plicial sets, Springer Lecture Notes in Math. 658 (1978), 80–150. [13]. A.K. Bousfield and V.K.A.M. Guggenheim, On PL de Rham theory and rational homo-

topy type, Mem. AMS 179 (1976). [14]. A.K. Bousfield and D.M. Kan, Homotopy limits completions and localizations, Springer

Lecture Notes in Math., Vol. 304 (2^{nd} corrected printing), Springer-Verlag, Berlin-Heidel-berg-New York, 1987.

[15]. K.S. Brown, Abstract homotopy theory and generalized sheaf cohomology, Trans. AMS

186 (1973), 419–458. [16]. K.S. Brown, Buildings, Springer-Verlag, Berlin-Heidelberg-New York, 1989. [17]. C. Cascuberta, Recent advances in unstable localization, The Hilton Symposium (1993),

CRM Proc. Lecture Notes 6, AMS, Providence, 1994, pp. 1–22. [18]. E.B. Curtis, Simplicial homotopy theory, Advances in Math. 6 (1971), 107–209. [19]. J. Dieudonné, Les débuts de la topologie algébrique, Expo. Math. 3 (1985), 347–357. [20]. A. Dold and D. Puppe, Homologie nicht-additiver Funktoren; Anwendungen, Ann. Inst.

Fourier (Grenoble) 11 (1961), 201–312. [21]. E. Dror, Pro-nilpotent representations of homotopy types, Proc. AMS 38 (1973), 657–660. [22]. E. Dror-Farjoun, Homotopy localization and v_{1}-periodic spaces, Algebraic Topology (San

Feliu de Guixols, 1990), Lecture Notes in Math. 1622, Springer-Verlag, Berlin, 1992, pp. 104–113.

[23]. E. Dror-Farjoun, Cellular spaces, null spaces and homotopy localization, Lecture Notes

in Math. 1622, Springer-Verlag, Berlin, 1996. [24]. W.G. Dwyer and D.M. Kan, Singular functors and realization functors, Ind. Math. 46

(1984), 147–153.

[25]. W.G. Dwyer and D.M. Kan, Homotopy theory and simplicial groupoids, Indag. Math. 46

(1984), 379–385. [26]. W.G. Dwyer and D.M. Kan, Equivalences between homotopy theories of diagrams, Alge-

braic Topology and Algebraic K-theory, W. Browder, ed., Annals of Math. Studies, Vol. 113, pp. 180–205, Princeton Univ. Press, Princeton, 1987.

[27]. W.G. Dwyer, D.M. Kan and J.H. Smith, Homotopy commutative diagrams and their

realizations, J. Pure Applied Albegra 57 (1989), 5–24. [28]. W.G. Dwyer, D.M. Kan and C.R. Stover, An E^{2} model category structure for pointed

simplicial spaces, Preprint. [29]. W.G. Dwyer, D.M. Kan and C.R. Stover, The bigraded homotopy groups π_{i,j}X of a

pointed simplicial space X, Preprint. [30]. W.G. Dwyer and J. Spalinski, Homotopy theories and model categories, Handbook of

Algebraic Topology, Elsevier, Amsterdam, 1995, pp. 73–126. [31]. S. Eilenberg, Singular homology theory, Ann. of Math. 45 (1944), 407–477. [32]. S. Eilenberg and J.A. Zilber, On semi-simplicial complexes and singular homology, Ann.

of Math. 51 (1950), 409–513. [33]. P. Gabriel and M. Zisman, Calculus of Fractions and Homotopy Theory, Springer-Verlag,

New York, 1967. [34]. H. Gillet and D. Grayson, The loop space of the Q-construction, Ill. J. Math. 31(4)

(1987), 574–597. [35]. P. Goerss, André-Quillen cohomology and the homotopy groups of mapping spaces: un-

derstanding the E_{2} term of the Bousfield-Kan spectral sequence, J. of Pure and Applied Algebra 63 (1990), 113–153.

[36]. P.G. Goerss, Homotopy fixed points for Galois groups, Preprint. [37]. P.G. Goerss, Comparing completions of a space at a prime, Preprint (1997). [38]. P.G. Goerss and J.F. Jardine, Localization theories for simplicial presheaves, Preprint

(1996). [39]. J.P.C. Greenlees and J.P. May, Derived functors of I-adic completion and local homology,

J. Algebra 149 (1992), 438–453. [40]. T. Gunnarsson, Abstract Homotopy Theory and Related Topics, Thesis, Chalmers Tek-

niska Högskola, Göteborg, 1978. [41]. J-C. Haussmann and D. Husemoller, Acyclic maps, Enseign. Math. 25 (1979), 53–75. [42]. P.S. Hirschhorn, Localization of Model Categories, http://www-math.mit.edu/∼psh. [43]. L. Illusie, Complexe cotangent et déformations I, Springer Lecture Notes in Math., Vol.

239, Springer-Verlag, Berlin-Heidelberg-New York, 1971. [44]. U. Jannsen, Continuous étale cohomology, Math. Ann. 280 (1988), 207–245. [45]. J.F. Jardine, Simplicial objects in a Grothendieck topos, Contemporary Math. 55 (1986),

193–239. [46]. J.F. Jardine, Simplicial presheaves, J. Pure Applied Algebra 47 (1987), 35–87. [47]. J.F. Jardine, Stable homotopy theory of simplicial presheaves, Can. J. Math. 39(3)

(1987), 733–747. [48]. J.F. Jardine, The homotopical foundations of algebraic K-theory, Contemporary Math.

83 (1989), 57–82. [49]. J.F. Jardine, Supercoherence, J. Pure Applied Algebra 75 (1991), 103–194. [50]. J.F. Jardine, Homotopy and homotopical algebra, Handbook of Algebra, Vol. 1, Elsevier,

Amsterdam, 1996, pp. 639–670. [51]. J.F. Jardine, Boolean localization, in practice, Doc. Math. 1 (1996), 245–275.

[52]. J.F. Jardine, Generalized Etale Cohomology Theories, Progress in Math., Vol. 146, Birk-

häuser, Boston-Basel-Berlin, 1997. [53]. A. Joyal, letter to A. Grothendieck, (1984). [54]. D.M. Kan, Semisimplicial spectra, Ill. J. Math. 7 (1963), 463–478. [55]. D.M. Kan, Abstract homotopy I, Proc. of Nat. Acad. of Sci. 41 (1955), 1092–1096. [56]. D.M. Kan, Abstract homotopy II, Proc. of Nat. Acad. of Sci. 42 (1956), 225–228. [57]. D.M. Kan, Abstract homotopy III, Proc. of Nat. Acad. of Sci. 42 (1956), 419–421. [58]. D.M. Kan, On c.s.s. complexes, Amer. J. Math. 79 (1957), 449–476. [59]. D.M. Kan, Adjoint functors, Trans. AMS 87 (1958), 294–329. [60]. D.M. Kan, A combinatorial definition of homotopy groups, Ann. of Math. 67 (1958),

288–312. [61]. K. Lamotke, Semisimpliziale algebraische Topologie, Springer-Verlag, Berlin-Heidelberg-

New York, 1968. [62]. J. Lannes, Sur les espaces fonctionnels dont la source est le classifiant d’un p-groupe

abélien élémentaire, I.H.E.S. Publications Mathématiques 75 (1992), 135–244. [63]. L.G. Lewis, J.P. May and M. Steinberger, Equivariant Stable Homotopy Theory, Springer

Lecture Notes in Math., Vol. 1213, Springer-Verlag, Berlin-Heidelberg-New York, 1986. [64]. S. Mac Lane, Homology, Third Corrected Printing, Springer-Verlag, New York Heidelberg

Berlin, 1975. [65]. S. Mac Lane, Natural associativity and commutativity, Rice University Studies 49 (1963),

28–46. [66]. S. Mac Lane, Categories for the Working Mathematician, Graduate Texts in Math., Vol.

5, Springer, Berlin-Heidelberg-New York, 1971. [67]. J.P. May, Simplicial Objects in Algebraic Topology, Van Nostrand, Princeton, 1967. [68]. J.P. May (with contributions by F. Quinn, N. Ray, and J. Tornehave), E_{∞} ring spaces

and E_{∞} ring spectra, Springer Lecture Notes in Math., Vol. 577, Springer-Verlag, Berlin-Heidelberg-New York, 1977.

[69]. J.P. May, Pairings of categories and spectra, J. Pure Applied Algebra 19 (1980), 299–346. [70]. D. McDuff and G. Segal, Homology fibrations and the group-completion theorem, Invent.

Math 31 (1976), 279–287. [71]. H. Miller, The Sullivan conjecture on maps from classifying spaces, Ann. of Math. (2)

120 (1984), 39–87. [72]. J.S. Milne, Étale Cohomology, Princeton University Press, Princeton, 1980. [73]. I. Moerdijk, Bisimplicial sets and the group-completion theorem, Algebraic K-Theory:

Connections with Geometry and Topology, NATO ASI Series C, Vol. 279, pp. 225–240, Kluwer, Dordrecht, 1989.

[74]. J.C. Moore, Semi-simplicial complexes and Postnikov systems, Symposium Internationale

de Topologica Algebraica, 1958, pp. 232–247. [75]. D. Quillen, Homotopical Algebra, Lecture Notes in Math., Vol. 43, Springer, Berlin-

Heidelberg- New York, 1967. [76]. D. Quillen, The geometric realization of a Kan fibration is a Serre fibration, Proc. AMS

19 (1968), 1499–1500. [77]. D. Quillen, Rational homotopy theory, Ann. of Math. 90 (1969), 205–295. [78]. D. Quillen, Higher algebraic K-theory I, Springer Lecture Notes in Math. 341 (1973),

85–147. [79]. D. Quillen, Homotopy properties of the poset of non-trivial p-subgroups of a group, Ad-

vances in Math. 28 (1978), 101–128.

[80]. C.L. Reedy, Homotopy theory of model categories, Preprint (1973). [81]. M. Ronan, Lectures on Buildings, Perspectives in Math., Vol. 7, Academic Press, Boston,

1989. [82]. H. Schubert, Categories, Springer, New York-Heidelberg-Berlin, 1972. [83]. G. Segal, Categories and cohomology theories, Topology 13 (1974), 293–312. [84]. J-P. Serre, Trees, Springer-Verlag, Berlin-Heidelberg-New York, 1980. [85]. D. Sullivan, Genetics of homotopy theory and the Adams conjecture, Ann. of Math. 100

(1974), 1–79. [86]. E.H. Spanier, Algebraic Topology, First corrected Springer edition, Springer-Verlag, New

York, 1981. [87]. M. Sweedler, Hopf Algebras, W.A. Benjamin, New York, 1969. [88]. R.M. Switzer, Algebraic Topology – Homotopy and Homology, Springer-Verlag, Berlin-

Heidelberg-New York, 1975. [89]. D.H. Van Osdol, Simplicial homotopy in an exact category, Amer. J. Math. 99 (1977),

1193–1204. [90]. R.M. Vogt, Homotopy limits and colimits, Math. Z. 134 (1973), 11–52. [91]. F. Waldhausen, Algebraic K-theory of spaces, Springer Lecture Notes in Math. 1126

(1985), 318–419.